OpenStudy (anonymous):

derive the value of whole root of1+sinx/1_sinx

OpenStudy (anonymous):

if you want the value, then you must supply a value for x

OpenStudy (anonymous):

simplification needed

OpenStudy (anonymous):

multiply top and bottom by conjugate

OpenStudy (anonymous):

it will simplify to something like sinx

OpenStudy (anonymous):

dammit that is right. I didn't think of that!

OpenStudy (anonymous):

or similar

OpenStudy (dumbcow):

what do you mean by root? like a square root or find the zero

OpenStudy (anonymous):

\[\sqrt{\frac{1+\sin(x)}{1-\sin(x)}}\]

OpenStudy (anonymous):

first ignore the sqrt

OpenStudy (anonymous):

\[\frac{1+\sin(x)}{1-\sin(x)} \times \frac{1+\sin(x)}{1+\sin(x)} \]

OpenStudy (dumbcow):

ahh ok then

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

should multiply by 1-sinx

OpenStudy (anonymous):

but its same idea

OpenStudy (dumbcow):

(1+sinx)/cosx

OpenStudy (dumbcow):

or secx + tanx

OpenStudy (anonymous):

you will get some 1-sin^2 , which can be changed to cos^2 , then the sqrt disappears when you include the sqrt

OpenStudy (dumbcow):

no elecengineer you were right first time, conjugate is 1+sinx

OpenStudy (anonymous):

nah but i remember certain times you dont multiply by the conjugate of the bottom, when you evaluate certain limits you dont , it all depends on whether you want the diff of squares on the top or bottom

OpenStudy (anonymous):

ah dont worry